### v2

# Advancements in Visual Language Models for Remote Sensing: Datasets, Capabilities, and Enhancement Techniques

Abstract
Recently, the remarkable success of ChatGPT has sparked a renewed wave of interest in artificial intelligence (AI), and the advancements in visual language models (VLMs) have pushed this enthusiasm to new heights. Differring from previous AI approaches that generally formulated different tasks as discriminative models, VLMs frame tasks as generative models and align language with visual information, enabling the handling of more challenging problems. The remote sensing (RS) field, a highly practical domain, has also embraced this new trend and introduced several VLM-based RS methods that have demonstrated promising performance and enormous potential. In this paper, we first review the fundamental theories related to VLM, then summarize the datasets constructed for VLMs in remote sensing and the various tasks they addressed. Finally, we categorize the improvement methods into three main parts according to the core components of VLMs and provide a detailed introduction and comparison of these methods. A project associated with this review has been created at https://github.com/taolijie11111/VLMs-in-RS-review.
Keywords Visual Language Models $\cdot$ Remote Sensing
1 Introduction
Remote sensing technology can help people obtain various observation data from a distance, and it is widely applied in fields such as disaster monitoring [1], urban planning [2], and agricultural management [3], surface observation [4], environmental monitoring [5], etc. Consequently, remote sensing image processing technology has been a key topic of focus for researchers both domestically and internationally. Over the past decade, with the development of artificial intelligence technology, significant breakthroughs have also been made in remote sensing image processing techniques. Specifically, various 3D convolutional neural networks (CNNs) [6] have been specifically designed for hyperspectral and multispectral image classification, as well as spectral reconstruction. U-Net [7] is utilized for semantic segmentation and land cover mapping in aerial imagery. A pyramid network has been proposed for multi-scale object detection in SAR images. The YOLO framework [8] is applied for small target detection in infrared remote sensing images. Besides, AI techniques are also applied in fields such as denoising, fusion, enhancement, and compression of remote sensing images, achieving remarkable results. In summary, AI techniques have significantly advanced the processing of remote sensing images, elevating the capabilities in these fields. However, most of these methods primarily focus on image processing, and the models they employ are discriminative models. There are inherent limitations to this type of technology, such as the inability to incorporate some human common sense, and the trained models can only perform a single vision task.
Recently, due to advancements in data availability, computational power, and the introduction of the transformer model [9], natural language processing (NLP) has made significant breakthroughs. By feeding tens of billions of text tokens and scaling the model to tens of billions of parameters, ChatGPT [10] achieved amazing performance in natural language understanding, language translation, text generation, question answering, etc. The impressive success of ChatGPT [10] has reignited enthusiasm for AI. Larger and more advanced models continue to emerge, showcasing unprecedented advancements in language understanding through their extensive knowledge and sophisticated reasoning abilities, demonstrating human-like thinking. Specifically, Google released the encoder-only model BERT [11], which pushed the GLUE [12] benchmark to 80.4% and achieved an accuracy of 86.7% on MultiNLI [13] and set a new record of 93.2 for the F1 score on the SQuAD v1.1 [14] question-answering test, surpassing human performance by 2.0 points. Meta released LLaMA [15], which contains a series of models for 7 billion parameters to 65 billion parameters and it can beat GPT-3 [16] with 1/10 parameters. The development of large language models reveals new possibilities for the advancement of artificial intelligence technology, particularly in generative AI. Compared to discriminative AI, generative AI models approach problem-solving in a more flexible and open manner, allowing for the modeling of a broader range of issues. The development of LLMs also offers new insights for addressing the challenges faced by discriminative models in the computer vision field.
The introduction of models such as CLIP [17] has effectively bridged the modality gap between images and language, creating opportunities for innovative integration. CLIP [17], with its innovative multimodal encoder-decoder structure utilizing BERT [11] as a text encoder and Vision Transformer (ViT) [18] / ResNet [19] as an image encoder, exemplifies the synergy between images and language vocabulary, leading to improved accuracy in classification and detection tasks through a deep understanding of the relationship between image content and textual vocabulary . GLIP [20] further enhances this consistency between language sentences and image content. Furthermore, the emergence of VLM has fundamentally addressed numerous constraints associated with traditional discriminative models in remote sensing image processing, thereby catalyzing the transition to AI 2.0. Generally speaking, VLM is a type of model designed to integrate and understand both visual and textual information, enabling tasks such as image captioning, visual question answering, and cross-modal retrieval by leveraging the relationships between images and language. Compared to previous discriminative models, VLM offers greater flexibility in modeling various AI tasks, allowing it to handle multiple tasks within a single framework. Additionally, its multimodal perceptual ability brings it closer to human perception. In the past two years, both LLMs and VLMs have developed rapidly, with various stronger models being introduced, such as LLaMA [15], LLaVA [21], and GPT-4 [22], among others. The performance of related tasks has also been continually improved.
Similar to previous discriminative AI technologies, which have significantly improved the field of remote sensing data processing, VLM—one of the core technologies in AI 2.0—also brings new opportunities for this domain. VLMs, by leveraging their multimodal capabilities, have demonstrated significant progress across various tasks within remote sensing, including geophysical classification [23, 24], object detection [25, 26], and scene understanding [27, 28]. By introducing LLaVA-1.5 [21] into the remote sensing field, Kuckreja et al. proposed GeoChat [25], which not only answers image-level queries but also accepts region inputs for region-specific dialogue [25]. Deng et al. designed ChangeChat, which utilizes multimodal instruction tuning to handle complex queries such as change captioning, category-specific quantification, and change localization [29]. Inspired by GPT-4 [22], Hu et al. developed a high-quality Remote Sensing Image Captioning dataset (RSICap) to facilitate the advancement of large VLMs in the remote sensing domain [30]. However, overall, the application of VLM technology in the field of remote sensing is still in its early stages. Current VLM-based remote sensing data processing technologies have already demonstrated significant potential, but there are still many issues that need to be addressed.
Contribution. In this paper, we present a comprehensive review of VLM-based remote sensing applications, offering a thorough delineation of the principal methodologies and the most recent advancements in VLMs, particularly within the context of remote sensing. By providing an in-depth introduction to the most representative and influential works in this domain, the survey enables researchers to quickly acquire a nuanced understanding of the developmental trajectory of these methodologies, facilitating the classification and evaluation of different approaches. Furthermore, the survey extends its contribution by presenting a detailed classification of enhancement techniques, focusing on key components such as vision encoders, text encoders, and the alignment between vision and language. This classification not only systematically illustrates the research progress made within this specialized field but also serves as a valuable reference point, supporting and inspiring further innovative research endeavors by scholars aiming to explore and advance the boundaries of VLM applications in remote sensing.
Survey Pipeline This survey is organised as follows. The section 2 presents an convice introduction to foundation models, including Transformer, Vision Transformer and the famious VLM—LLaVA. The section 3 outline the datasets used in VLM for RS, which can be further divided manually into collected data, data combining with existing dataset and data enhanced by VLMs and LLMs. The section 4 describe the capabilities of VLMs, specifically introduce the pure visual tasks and visual language tasks. In Section 5, current VLM-based remote sensing data processing methods are introduced according to three major improvement directions, with a comparison and analysis of different approaches. Finally, several promising future directions are discussed in Section 6.
Figure 1: The illustration of Transformer [9]
2 Foundation Models
2.1 Transformer
[…]
$Attention(Q,K,V)=Softmax\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
where $d_{k}$ is a scaling factor, typically set to the dimension of the input feature head. Q, K, and V are variations of the input data, usually generated by a linear layer. For instance, $Q=W_{q}(x)$, $K=W_{k}(x)$, $V=W_{v}(x)$, where $x$ donates the input feature and $W_{q}$$W_{k}$ and $W_{v}$ are learnable parameters of the linear layers. When Q,K and V are variations of the same input feature, it is referred to as self-attention. If Q varies from $x_{1}$ while K and V vary from $x_{2}$ (with $x_{1}$ and $x_{2}$ coming from different sources), it is called cross-attention.
The Transformer model offers several advantages over traditional RNNs and LSTMs. Firstly, it utilizes self-attention mechanisms, allowing it to process input sequences in parallel rather than sequentially, which significantly speeds up training. Secondly, the ability to capture long-range dependencies without the vanishing gradient problem enhances its effectiveness in understanding context. This efficiency and scalability enable the development of large language models (LLMs), as Transformers can handle vast amounts of data and complex patterns more effectively than RNNs and LSTMs, paving the way for advancements in natural language processing tasks.
2.2 Vision Transformer
Figure 2: The illustration of Vision Transformer [18]
Inspired by the success and popular of transformer used in NLP, the Vision Transformer (ViT) model was proposed later and tailored for computer vision field. As shown in Figure 2, ViT is composed of three components, that is Linear Projection of Flattened Patches(Embedding layer), Transformer Encoder and MLP Head. In summary, to enable NLP models to process images, the images are serialized by cropping them into uniform-sized patches. These patches are then transformed into features using a projection layer. To ensure the model is sensitive to the positional information of the image patches, position embeddings are added to the corresponding patch embeddings. Currently, there are four types of position embedding methods, including 1) absolute position embedding; 2) relative position embedding (RPE) [31]; 3) conditional position embedding (CPE) [32]; 4) rotary position embedding (RoPE) [33]. The ViT adopts the first one.
Compared to convolutional networks, ViT has a significant advantage in scaling up ability. Its performance on image processing does not saturate quickly as the model size increases, allowing for better utilization of larger models and enhanced capabilities in handling more complex tasks. This characteristic has sparked a surge of research interest in the ViT architecture, leading to the continuous emergence of various improved versions of ViT models, such as SwinTransformer [34], Deit [35], etc. So far, the model paradigms in both the NLP and CV fields have been unified, which has laid the foundation for the emergence of VLM.
2.3 Vision Language Model
Figure 3: The illustration of LLaVA [21]
Figure 4: Ten representative tasks of VLM [36]
The introduction of the Transformer has significantly advanced language processing, while the emergence of ViT has unified the processing paradigms in NLP and CV. CLIP [17] further broke down the modal barriers between language and images. Building on these foundations, VLMs have been proposed by merging ViT and LLMs. Figure 3 shows the architecture of LLaVA [21], a very representive VLM model. From Figure 3, we can see that the core idea of LLaVA is to combine visual tokens with text tokens and feed them into a large language model (LLM) to generate text-based answers. Pre-trained CLIP visual encoder is adopted here to extract visual tokens, and a trainable projection layer is build to convert visual tokens to language embedding tokens.
Although the merging approach is straightforward, the resulting VLM can handle a wider range of tasks and interact with users, enabling conversations based on the content of images. Figure 4 illustrates ten representative tasks of VLM. From the results, it is evident that the model can not only perceive the content of images but also possesses a certain degree of reasoning ability, as demonstrated in the example shown in the bottom left corner of the Figure 4.
3 Datasets
Figure 5: Comparison between traditional datasets and VLM dataset in remote sensing.
Whether it is the previous discriminative models or the current generative models, they all fall under data-driven models. Therefore, VLM-based remote sensing image processing methods also require corresponding training data as support. Figure 5 compares the datasets used in previous discriminative models with those used in current VLMs. From the figure, we can observe that the most significant improvement is that VLM datasets include both image and language annotations.
In the remote sensing field, existing VLM datasets can be broadly categorized into three types based on their production approaches:
  • Manual Datasets, which are completely manually annotated by humans based on specific task requirements.
  • Combining datasets, which are build by combing several existing datasets and adding part new language annotations.
  • Automatically annotated datasets. This type of data construction involves minimal human participation and relies on various multimodal models for filtering and annotating image data.
In this section, we will delve into the construction methods and applicable task domains of existing remote sensing VLMs datasets, offering insights into how these datasets support extensive applications and in-depth research within the field. The statistical infromation of current existing datasets are summarized in Table 1
Table 1: The Statistical Information of Datasets. K=Thousands, M=Milion, SC=Scene Classification, VQA=Visual Question Answer, VG=Visual Grounding, RRSIS=Referring Remote Sensing Image Segmentation, IC=Image Captioning.
Method Dataset Source Image Image Size Task
Manual HallusionBench [37] - 346 - VQA
RSICap [30] DOTA 2585 512 IC
CRSVQA [38] AID 4639 600 VQA
Combining with Existing Datasets SATIN [39] Million-AID,WHU-RS19, SAT-4,AID1 $\approx$775K - SC
GeoPile [40] NAIP,RSD46-WHU, MLRSNet,RESISC45, PatternNet 600K - -
SatlasPretrain [41] UCM,BigEarthNet, AID,Million-AID, RESISC45,FMoW, DOTA,iSAID 856K 512 -
RSVGD [42] DIOR 17402 800 VG
RefsegRS [43] SkyScapes 4420 512 RRSIS
SkyEye-968K [44] RSICD,RSITMD, RSIVQA,RSVG1 968K - -
MMRS-1M [45] AID,RSIVQA, Syndney-Captions1 1M - -
RSSA [46] DOTA-v2,FAIR1M 44K 512 -
FineGrip [47] MAR20 2649 - -
RRSIS-D [48] - 17402 800 RRSIS
RingMo [49] Gaofen,GeoEye, WorldView,QuickBird1 2096640 448 -
GRAFT [50] NAIP,Sentinel-2 - - -
SkySense [51] WorldView,Sentinel-1, Sentinel-2 21.5M - -
EarthVQA [52] LoveDA,WorldView 6000 - VQA
GeoSense [53] Sentinel-2,Gaofen, Landsat,QuickBird $\approx$9M 224 -
Enhenced by VLMs and LLMs RS5M [26] LAION2B-en, LAION400M, LAIONCOCO1 5M - -
SkyScript [27] Google Earth Engine, OpenStreetMap 2.6M - -
LHRS-Align [54] Google Earth Engine, OpenStreetMap 1.15M - -
GeoChat [25] SAMRS,NWPU-RESISC-45, LRBEN,Floodnet 318k - -
GeoReasoner [55] Google Street View, OpenStreetMap 70K+ - StreetView
HqDC-1.4M [46] Million-AID, CrowdAI, fMoW,CVUSA, CVACT,LoveDA $\approx$1.4M 512 -
ChatEarthNet [56] Sentinel-2,WorldCover 163488 256 -
VRSBench [57] DOTA-v2,DIOR 29614 512 -
FIT-RS [58] STAR 1800.8K 512 -
1 Only a subset of common datasets are enumerated.
3.1 Manual Datasets
Manually curated datasets, though limited in size, are typically of high quality and are specifically designed for tasks. There are three high-quality datasets made by hand, which are meticulously crafted to address specific challenges within the remote sensing domain.
HallusionBench [37], the first benchmark designed to examine visual illusion and knowledge hallucination of large visual language models(LVLMs) and analyze the potential failure modes based on each hand-crafted example pair, consists of 455 visual-question control pairs, including 346 different figures and a total of 1129 questions on diverse topics. RSICap [30], RSIEval comprises five experts annotated captions with rich and high-quality information, thus the model achieves satisfactory performance with fine-tuning. CRSVQA [38] employs annotators from different backgrounds, including geography, surveying, mapping, to write complex questions that should reason the relationship between the objects or a concise question related to the image content. This can effectively prevent the problems of data redundancy and language bias in existing RSVQA datasets [59].
3.2 Combining Datasets
As shown in Table 1, some datasets are built from scratch, whose sources are usually from Sentinel-2, Gaofen, Google Earth Engine and so on. But most datasets created by merging existing remote sensing datasets, which facilitate multi-task models by integrating various data types from traditional RS domain-specific datasets, greatly eases labour-intensive compared to manual production and starting from scratch. The most utilized datasets are AID [60], NWPU-RESISC45 [61], UCM [62] for Scene Classification, FAIR1M [63], DIOR [64] for Object Detection, iSAID [65] for Semantic Segmentation, LEVIR-CD [66] for Change Detection, more details about those datasets could be found in section 4. Besides, the well-known benchmark dataset Million-AID [67] is frequently included.
3.3 Automatically Annoteted Datasets
The explosive development of VLMs and LLMs has provided efficient methods for constructing large-scale remote-sensing visual language datasets. This category, which is rapidly becoming the mainstream, leverages the power of VLMs and LLMs to construct large-scale, high-quality datasets.
RS5M [26] employs the general multimodal model BLIP2 [68] to generate image captions and then uses CLIP [17] to select the top 5 highest-scoring captions. SkyScript [27] matches Google Images with the OpenStreetMap (OSM) database and selects relevant image attribute values from OSM to concatenate them with CLIP [17] as image captions. LHRS-Align [54] uses a similar approach to extract multiple attribute values from OSM and then summarizes them using Vicuna-13B [69] to produce long and fluid captions. GeoChat [25] specifically provide system instructions as prompts that ask Vicuna [69] to generate multi-round question and answer pairs. GeoReasoner [55] obtaines image-text data pairs for geolocation from the communities of the two games, utilizes BERT [11] to clean and filter text that lacked geolocation-specific information, and then uses CLIP [17] to align the image and text to construct a highly locatable Street View dataset. HqDC-1.4M [46] utilized the “gemini-1.0-pro-vision" [70] to generate descriptions for images from multiple public RS datasets, thereby obtaining a dataset of image-text pairs to serve as the pretraining data for RSVLMs. ChatEarthNet [56] carefully designed prompts that embed semantic information from the land cover maps to make ChatGPT-3.5 and ChatGPT-4v achieve the best results on satellite imagery for the caption generation. VRSBench [57] did a similar job and spent thousands of hours manually verifying the generated annotations. FIT-RS [58] used the efficient TinyLLaVA-3.1B [71] to swiftly generate concise background descriptions for numerous RSIs, then combined them as the prompt to obtain detailed and fluent sentences using GPT-4/GPT-3.5. Examples, which utilize advanced models like BLIP2[68] and CLIP [17] to generate detailed image-text pairs, supporting complex applications such as disaster monitoring, urban planning, and environmental analysis.
3.4 Summary
The existing three types of datasets each have their own advantages and disadvantages, as detailed below:
  • Manually annotated datasets generally have higher quality and are most closely aligned with specific problems. But, these datasets typically consisting of only a few hundred or thousand images, are limited in scale compared to combined datasets and VLM datasets, which can reach tens of millions of images. The size of a dataset is crucial for model training, as larger and richer datasets yield more accurate and versatile models. Manually annotated datasets can be used to fine-tune large models, but they struggle to support training large models from scratch.
  • Combining datasets. By merging existing datasets, large-scale datasets can be constructed at a low cost. This type of data can easily reach millions of entries. However, datasets created in this manner typically have lower annotation quality compared to manually annotated datasets and may not fully align with specific tasks. Nonetheless, they can be utilized for model pre-training.
  • Automatically annotated datasets. Mainstream VLM dataset production methods leverage state-of-the-art LLMs and VLMs to generate high-quality captions or annotations for large-scale remote sensing images. This approach produces flexible image-text pairs, including long sentences, comments, and local descriptions, without requiring human supervision, thus meeting the needs of various tasks. Researchers may need to craft prompts and commands to guide LLMs and VLMs in generating the desired texts.
Overall, automatically annotated datasets leverage the advantages of pre-trained large models, resulting in datasets that are both quantitatively and qualitatively substantial. As a result, this type of dataset has become quite mainstream in the field of remote sensing.
Figure 6: Examples of tasks handled by VLMs in remote sensing field
4 Capabilities
In various tasks within the field of remote sensing, VLMs have demonstrated powerful capabilities. These capabilities can be broadly classified into two categories: pure visual tasks and visual language tasks. Pure visual tasks focus on analyzing remote sensing images to extract meaningful information about the Earth’s surface. Visual language tasks, which combine natural language processing with visual data analysis, have opened new avenues for remote sensing applications. The zero-shot and Few-shot tasks are SC, OD, SS and IR, which are described in more detail below. Figure 6 illustrates each capability with a simple example, particularly highlighting two emerging tasks along with their source tasks.
4.1 Pure visual
  • Scene Classification(SC) RS scene classification entails the classification of satellite or aerial images into distinct land-cover or land-use categories with the objective of deriving valuable insights about the Earth’s surface. It offers valuable information regarding the spatial distribution and temporal changes in various land-cover categories, such as forests, agricultural fields, water bodies, urban areas, and natural landscapes. The three most commonly used datasets in this field are AID [60],NWPU-RESISC45 [61] and UCM [62], which contain 10,000, 31,500, 2,100 images, and 30,45,21 categories, respectively.
  • Object Detection(OD) RSOD aims to determine whether or not objects of interest exist in a given RSI and to return the category and position of each predicted object. In this domain, the most widely used datasets are DOTA [72] and DIOR [64]. DOTA [72] contains 2806 aerial images which are annotated by experts in aerial image interpretation, with respect to 15 common object categories. DIOR [64] contains 23463 images and 192472 instances, covering 20 object classes, much more details than DOTA [72].
  • Semantic Segmentation(SS) Semantic Segmentation is a computer vision task in which the goal is to categorize each pixel in an image into a class or object. For RS images, it plays an irreplaceable role in disaster assessment, crop yield estimation, and land change monitoring. The most widely used datasets are ISPRS Vaihingen and Potsdam, which contain 33 true orthophoto (TOP) images and 38 image tiles with a spatial resolution of 5 cm, respectively. Another widely used dataset iSAID [65] contains 2806 aerial images, which mainly collected from Google Earth.
  • Change Detection(CD) Change detection in remote sensing refers to the process of identifying differences in the structure of objects and phenomena on Earth surface by analysing two or more images taken at different times. It plays an important role in urban expansion, deforestation, and damage assessment. LEVIR-CD [66], AICD [73] and the Google Data Set focus on changes of buildings. Other types of changes such as roads,groundwork are commonly included in CD datasets.
  • Object Counting(OC) RSOC aims to automatically estimate the number of object instances in a RSI. It plays an important role in many areas, such as urban planning, disaster response and assessment, environment control and mapping. RemoteCount [24] is a manual dataset for object counting in remote sensing imagery consisted of 947 image-text pairs and 13 categories, which are mainly selected from the validation set of the DOTA dataset.
4.2 Combine with natural language
  • Image Retrieval(IR) IR from RS aims at retrieving the interested RS images from the massive RS image repositories. With the launch of more and more Earth observation satellites and the emergence of large-scale remote sensing datasets, RS image retrieval can prepare auxiliary data or narrow the search space for a large number of RS image processing tasks.
  • Visual Question Answer(VQA) VQA is a task that seeks to provide answers to free-form and open-ended questions about a given image. As the questions can be unconstrained, a VQA model applied to remote sensing data could serve as a generic solution to classical problems but also very specific tasks involving relations between objects of different nature. VQA was first proposed by Anto et al. [74] in 2015, and then made it greatly developed in the RS domain by RSVQA [59] in 2020. It contributed a RSVQA framework and two VQA datasets, RSVQA-LR and RSVQA-HR, which wildly utilized in further works.
  • Image Captioning(IC) IC aims to generate natural language descriptions that summarizes the content of an image. It requires representing the semantic relationships among objects and generating an exact and descriptive sentence, so that it’s more complex and challenge than image detection, classification, and segmentation tasks. In this domain, the commonly used datasets are Syndney-Captions [75], RSICD [76], NWPU-Captions [77], and UCM-Captions [75]. Recently, CapERA [78] provids UAV video with diverse textual descriptions to advance visual–language-understanding tasks.
  • Visual Grounding(VG) RSVG aims to localize the referred objects in remote sensing images with the guidance of natural language. It was first introduced in [79] in 2022. In 2023, Yang [42] et al. not only build the new large-scale benchmark of RSVG based on detection in DIOR dataset, termed RSVGD, but design a novel transformer-based MGVLF module to address the problems of scale variation and cluttered background of RS images.
  • Remote Sensing Image Change Captioning(RSICC) A new task aiming at generating human-like language descriptions for the land cover changes in multitemporal RS images. It is a combine of IC and CD task, offering important application prospects in damage assessment, environmental protection, and land planning. Chenyang et al. [80] firstly introduced the CC task into the RS domain in 2022 and proposed a large-scale dataset LEVIR-CC containing 10077 pairs of bitemporal RS images and 50385 sentences describing the differences between the images. In 2023, they propose a pure Transformer-based model [81] to further improve the performance of the RSICC task.
  • Referring Remote Sensing Image Segmentation(RRSIS) RRSIS is to provide a pixel-level mask of desired objects based on the content of given remote sensing images and natural language expressions. It was firstly proposed by Zhenghang et al. [43] in 2024, who created a new dataset, called RefSegRS, for this task. In the same year, Sihan et al. [48] curated an expansive dataset comprising 17,402 image-caption-mask triplets, called RRSIS-D, and propose a series of models to meet the ubiquitous rotational phenomena in RRSIS.
The invention of Visual Language Models (VLMs), has greatly changed the remote sensing field which mainly focus on geospatial analysis. With the development and usage of various sensors, conventional vision models can not handle such complex data type, while VLMs technique could enhance the accuracy and efficiency of remote sensing analysis. Specifically, the combination of visual and linguistic data within VLMs allow more subtle and sophisticated interpretation of complex remote sensing datasets. By utilizing their multimodal capabilities, VLMs have showed significant progress across various tasks within remote sensing, including object detection, geophysical classification and scene understanding. Meanwhile, it provides more accurate and smart solutions for practical applications such as disaster monitoring, urban planning, and agricultural management.
5 Recent Advances in Visual Language Models for Remote Sensing
Figure 7: Fundamental architecture of contrastive and conversational visual language models.
Currently, there are mainly two types of models that integrate vision and language: one is based on contrastive learning structures, such as the CLIP series, and the other is based on large language models that fuse visual features, like LLaVA. As shown in Figure 7. Accordingly, the current improvements can also be categorized into two major types: enhancements within the contrastive learning framework and improvements within the conversational framework.
5.1 Advancements in Contrastive Vision-Language Models
Table 2: Architecture of contrastive VLMs, SC=Scene Classification, VG=Visual Grounding, IC=Image Captioning, CD=Change Detection, IR=Image Retrieval. For each model, we indicate the LLM used in its best configuration as shown in the original paper.
Model Image Encoder Text Encoder Capabilitie
RemoteCLIP [24] ViT-B (CLIP) Transformer SC,IC,IR,OC
CRSR [82] ViT-B (CLIP) - IC
ProGEO [83] ViT-B (CLIP) BERT VG
GRAFT [50] ViT-B (CLIP) - SC,VQA,,IR
GeoRSCLIP [26] ViT-H (CLIP) - SC,IR
ChangeCLIP [84] CLIP - CD
APPLeNet [85] CLIP - SC,IC
MGVLF [42] ResNet-50 BERT VG
Contrastive VLMs primarily consist of an image encoder and a text encoder, with the core challenge being the alignment of features extracted by both encoders. The image encoder, utilizing a convolutional neural network (CNN) or Vision Transformer (ViT), encodes images into feature vectors. Concurrently, the text encoder processes textual descriptions into corresponding feature vectors. The model is trained to maximize the similarity between positive image-text pairs while minimizing the similarity between negative pairs. CLIP is a pioneering model in vision-language pre-training, employing InfoNCE loss to align text and image features in a shared space. Currently, in the field of remote sensing, contrastive VLMs focus on transferring the image-text alignment capabilities established in RGB images and text to remote sensing images and text, which are then utilized to address various downstream tasks, such as change detection and remote sensing image caption,
By converting annotations from 10 object detection datasets, 4 remote sensing semantic segmentation datasets, and 3 remote sensing image-text datasets using box-to-caption and mask-to-box conversions, Liu et al. created a combined image-caption dataset and fine-tuned the CLIP model on this newly constructed dataset through continual learning [24]. Similar to [24], GeoRSCLIP builds a remote sensing dataset consisting of 5 million image-text pairs by filtering existing image-text pair datasets and generating captions for image-only datasets. It then fine-tunes CLIP on this proposed dataset [26]. CRSR [82] incorporates a Transformer Mapper network, utilizing an attention mechanism and learnable semantic predictive queries to provide a comprehensive representation of image features. This approach effectively analyzes and understands more semantic details present in remote sensing imagery (RSI). ProGEO proposes adding more descriptions to a geographic image-based pretrained CLIP, using these text descriptions to enhance the image encoder by aligning features from the image encoder with the features of the descriptions [83]. This approach enables the image encoder to generate features with more details, leading to improvements in visual geo-localization. CRAFT propose to train a satellite image model to align its embedding with the CLIP embedding of co-located internet images, which sidestep the need for texual annotations for training remote-sensing VLMs [50]. ChangeCLIP [84] is the first to apply a multimodal vision-language scheme to Remote Sensing Change Detection (RSCD) tasks. It introduces a Differentiable Feature Calculation (DFC) layer, enabling the network to focus on learning the changed features of bitemporal remote sensing images rather than semantic categories. APPLeNet design a injection module to fuse image features from different level and interact with text to generate better visual tokens [85].
In addition, S-CLIP [86] and RS-CLIP [87] enhance the performance of CLIP in remote sensing by improving the pseudo-label strategy. The former divides pseudo-labels into title-level and keyword-level categories, combining their advantages to ensure the model’s generalization ability. RS-CLIP automatically generates pseudo-labels from unlabeled data and employs a curriculum learning strategy to gradually select more pseudo-samples for model fine-tuning.
Figure 8: Enhancement Techniques
5.2 Advancements in Conversational Vision-Language Models
Currently, conversational VLMs are particularly noteworthy for their ability to address complex remote sensing tasks by transforming intricate visual features into a format that LLMs can comprehend. By integrating advanced image encoders that adeptly capture subtle variations in remote sensing imagery with text encoders optimized for efficiently processing language instructions, these models demonstrate remarkable performance, particularly in few-shot learning scenarios. The alignment layer is pivotal in establishing connections between remote sensing imagery and LLMs. Techniques such as utilizing MLPs for feature projection and incorporating learnable query embeddings for task-specific image feature extraction have proven effective in elevating the overall performance of VLMs.
Leveraging the existing API of pre-trained LLMs, conversational VLMs extract features from image data using a visual encoder. These features are then transformed into a format that LLMs can interpret, enabling the completion of various visual language tasks. To enhance task performance, researchers meticulously designed each component, which is described in detail below. Figure 8 clearly shows the model details of conversational VLMs.
Image Encoder Table 3 presents the visual encoders commonly used in current VLM implementations. Most studies utilizes the standard CLIP image encoder [17], but RSGPT [30] and SkyEyeGPT [44] have explored other variants, EVA-CLIP encoder [88], which is trained with improved training techniques.
Additionally, EarthGPT[45] adopt the vision transformer (ViT) as the image encoder to extract multi-layer features, which is beneficial to capture subtle differences in RS images and provides a more comprehensive understanding of images. Then the CNN backbone is designed as the image encoder to integrate multi-scale local details into visual representation. GeoChat [25], in addition to using CLIP-ViT(L-14) as the visual backbone network, interpolates the positional encoding in the transformer-based CLIP model with high-resolution 512x512 to understand details presented in remote sensing imagery (e.g., small objects and object details).
Table 3: Architecture of conversational VLMs, $\ast$: frozen, $\blacktriangle$: fine-tuning, SC=Scene Classification, VQA=Visual Question Answer, VG=Visual Grounding, IC=Image Captioning, OD=Object Detection, IR=Image Retrieval, RSICC=Remote Sensing Image Change Captioning. For each model, we indicate the LLM used in its best configuration as shown in the original paper.
Model Image Encoder Text Encoder Capabilitie
RS-LLaVA [89] ViT-L(CLIP) Vicuna-v1.5-13B IC,VQA
H2RSVLM [46] ViT-L (CLIP) Vicuna-v1.5-7B SC,VQA,VG
SkySenseGPT [58] ViT-L (CLIP) Vicuna-v1.5-7B SC,IC,VQA,OD
GeoChat [25] ViT-L (CLIP) Vicuna-v1.5-7B SC,VQA,VGRegion Caption
RSGPT [30] ViT-G (EVA) Vicuna-v1.5-7B IC,VQA
SkyEyeGPT [44] ViT-G(EVA) LLaMA2-7B IC,VQA,VG
RS-CapRet [90] ViT-L(CLIP) LLaMA2-7B IC,IR
LHRS-Bot [54] ViT-L (CLIP) LLaMA2-7B SC,VQA,VG
EarthGPT [45] ViT-L (DINOv2) ConvNeXt-L (CLIP) LLaMA2-7B SC,VQA,IC,VG,OD
Liu et al. [81] ViT-B(CLIP) GPT-2 RSICC
Text Encoder Training large language models (LLMs) from scratch is highly costly. Therefore, conversational visual language models (VLMs) utilize pre-trained LLMs as text encoders. These LLMs process visual elements extracted by the visual encoder and converted by middleware, along with the researcher’s questions or language instructions, to generate the desired results. Currently, the academic community is focusing on the LLaMA [21] and Vicuna [69] series, which serve as text encoders for most conversational LLMs. Table 3 provides a summary of these models. Common approaches include directly using LLMs or efficiently fine-tuning them with Low-Rank Adaptation (LoRA) [91]. Liu et al. [81] designed a complex image classifier and proposed a multi-prompt learning strategy to generate multiple learnable prompt embeddings as prefixes for LLMs to enhance their utilization for the RSICC task. RS-LLaVA [89] and RSGPT [30] demonstrated the capabilities of the proposed architecture using two different LLMs (vicuna-7B and vicuna-13B). The results indicated that fine-tuning larger language models leads to higher performance but at the cost of longer training times and higher computational resource requirements.
Vision-Language Alighment In the architecture of conversational Vision-Language Models (VLMs), the middle connection layer plays a crucial role. It acts as a bridge between remote sensing images and Large Language Models (LLMs), projecting image information into a space that LLMs can comprehend, enabling them to understand complex remote sensing (RS) tasks. A common approach to bridge the modal gap is to use Multi-Layer Perceptrons (MLPs), which employ single linear layer [44, 45] or double layers [46, 58, 89] to project visual tokens and align feature dimensions with word embeddings. Additionally, some works focus on designing connectors with intricate structures, offering valuable insights for future research. SkyEyeGPT [44], to balance the important visual information processing of high-resolution remote sensing images by LLMs and the high resource requirements, uses the method by Chen et al. [92] to directly concatenate 4 adjacent visual tokens, reducing the number of patches by 4 times.
One reason current general VLMs perform poorly in the remote sensing field is that the visual encoder of LLMs usually uses CLIP, but CLIP’s ViT lacks spatial awareness, affecting VLM’s spatial understanding ability [93]. To address this issue, H2RSVLM [46] and SkySenseGPT [58] enhance the visual encoder’s perceptual ability at the source by creating higher quality, large-scale, and detailed image-text pairs in remote sensing datasets. EarthGPT [45] expands the input images to include SAR and infrared images. Furthermore, some research endeavors opt for designing elaborate visual language connection structures to enhance the multimodal perception capabilities of VLM. GeoChat [25] constructs bounding boxes for text representations to express their spatial coordinates, enabling it to perform object grounding or accept region inputs, thus having reasoning capabilities for remote sensing images. RSGPT [30], under the guidance of InstructBLIP [94], inserts an annotation-aware query transformer (Q-Former) to enhance the alignment of visual and textual features. Additionally, the annotation-aware mechanism interacts with query embeddings through self-attention layers, allowing RSGPT to extract comprehensive visual features crucial for RS visual language tasks. LHRS-Bot [54] introduces a set of learnable queries for each level of image features. These queries aim to summarize the semantic information of each level of visual features through a series of stacked cross-attention and MLP layers.
Table 4: Summary of Other VLMs in section 5.3.
Model Enhancement Technique
Txt2Img [95] Employ modern Hopfield layers for hierarchical prototype learning of text and image embeddings.
CPSeg [96] Utilize "chain of thought" process that leverages textual information associated with the image.
SHRNet [97] Utilize a hash-based architecture to extract multi-scale image features.
MGeo [98] Introduce the geographic environment (GC) extracting multi-modal correlations for accurate query-POI matching.
GeoCLIP [99] Via an image-to-GPS retrieval approach by explicitly aligning image features with corresponding GPS locations.
SpectralGPT [100] Provide a foundational model for the spectral data based on the MAE[101] architecture.
TEMO [102] Capture long-term dependencies from description sentences of varying lengths, generating text modal features for each category.
5.3 Others
In addition to the aforementioned approaches, as shown in Table 4, several other advancements have been made to address domain-specific tasks in remote sensing. Txt2Img [95] employs modern Hopfield layers for hierarchical prototype learning of text and image embeddings, achieving notable performance in generating text descriptions for remote sensing images. CPSeg [96] enhances image segmentation by integrating a novel "chain of thought" process that leverages textual information associated with the image. SHRNet [97] utilizes a hash-based architecture to extract multi-scale image features, while QRN improves alignment efficiency and enhances the visual-spatial reasoning ability of remote sensing visual question answering (VQA) systems. MGeo [98], a multi-modal geographic language model, introduces the geographic environment (GC) as a new modality, effectively extracting multi-modal correlations for accurate query-POI matching. GeoCLIP [99], with its independently designed Location Encoder, is the first to tackle worldwide geo-localization via an image-to-GPS retrieval approach by explicitly aligning image features with corresponding GPS locations. SpectralGPT [100], based on the MAE[101] architecture, provides a foundational model for the spectral data domain, supporting various downstream tasks. TEMO [102] enhances the word embedding component of VLMs, capturing long-term dependencies from description sentences of varying lengths, generating text modal features for each category, and employing a masking strategy to prevent overfitting. This method is particularly effective for few-shot object detection (FSOD) in text-modal knowledge-oriented remote sensing imagery.
5.4 Performance Comparison
In this section, we present a comparative analysis of various Visual-Language Model (VLM) pre-training methods, focusing on their performance across common datasets. It is important to note that we have prioritized performance evaluations on widely recognized datasets, which has resulted in the exclusion of methods that are tailored for specific datasets.
Table  5 summarizes the evaluation results for five prevalent image classification tasks, highlighting the best-performing VLMs. Due to the differing implementations of VLM pre-training methods, it is crucial to identify the standout models. Among those utilizing a contrastive approach, RemoteCLIP [24] emerges as the top performer, while the conversational approach is epitomized by SkySenseGPT [58], which consistently demonstrates superior performance. Overall, our findings suggest that models employing conversational techniques tend to outperform their contrastive counterparts.
Table 5: Performance of VLM pre-training methods on common remote sensing datasets, the best comprehensive performance was deepened separately. $\ast$: few-shot prediction.
Category Model AID [60] RSVGD [42] NWPU-RESISC45 [61] WHU-RS19 [103] EuroSAT [104]
Contrastive RemoteCLIP [24] 91.30 - 79.84 96.12 60.21
GeoRSCLIP [26] 76.33 - 73.83 - 67.47
APPLeNet [85] - - 72.73 - -
GRAFT [50] - - - - 63.76
Conversational SkySenseGPT [58] 92.25 - - 97.02 -
LHRS-Bot [54] 91.26 88.10 - 93.17 51.40
H2RSVLM [46] 89.33 48.04 - 97.00 -
GeoChat [25] 72.03 - - - -
SkyEyeGPT [44] - 88.59 - - -
EarthGPT [45] - 81.54 93.84 - -
Others RSVG [42] - 78.41 - - -
Table 6: Performance of VLM pre-training methods on visual question answer task, the best comprehensive performance was deepened separately. $\ast$: few-shot prediction.
Model RSVQA-LR [59] RSVQA-HR [59] RSIVQA [105]
SkySenseGPT [58] 92.69 76.64 -
RSGPT [30] 92.29 92.00 -
GeoChat [25] 90.70 72.30 -
LHRS-Bot [54] 89.19 92.55 -
H2RSVLM [46] 89.12 74.35 -
RS-LLaVA [89] 88.56 - -
SkyEyeGPT [44] 88.23 86.87 -
EarthGPT [45] - 72.06 -
SHRNet [97] 85.85 85.39 84.46
MQVQA [38] - - 82.18
Table 7: Performance of VLM pre-training methods on image caption task, the best comprehensive performance was deepened separately.
Dataset Model BLEU1 BLEU2 BLEU3 BLEU4 METEOR ROUGEH CIDEr
UCM- Captions [62] CRSR [82] 90.60 85.61 81.22 76.81 49.56 85.86 380.69
SkyEyeGPT [44] 90.71 85.69 81.56 78.41 46.24 79.49 236.75
RS-LLaVA [89] 90.00 84.88 80.30 76.03 49.21 85.78 355.61
RSGPT [30] 86.12 79.14 72.31 65.74 42.21 78.34 333.23
RS-CapRet [90] 84.30 77.90 72.20 67.00 47.20 81.70 354.80
Syndney- Captions [75] CRSR [82] 79.94 74.4 69.87 66.02 41.50 74.88 289.00
SkyEyeGPT [44] 91.85 85.64 80.88 77.40 46.62 77.74 181.06
RSGPT [30] 82.26 75.28 68.57 62.23 41.37 74.77 273.08
RS-CapRet [90] 78.70 70.00 62.80 56.40 38.80 70.70 239.20
RSICD [76] CRSR [82] 81.92 71.71 63.07 55.74 40.15 71.34 306.87
SkyEyeGPT [44] 87.33 77.70 68.90 61.99 36.23 63.54 89.37
RSGPT [30] 70.32 54.23 44.02 36.83 30.1 53.34 102.94
RS-CapRet [90] 72.00 59.90 50.60 43.30 37.00 63.30 250.20
NWPU- Captions [77] RS-CapRet [90] 87.10 78.70 71.70 65.60 43.60 77.60 192.90
EarthGPT [45] 87.10 78.70 71.60 65.50 44.50 78.20 192.60
In Tables  6 and  7, we delve deeper into two frequently encountered visual-language tasks: Visual Question Answering (VQA) and Image Classification (IC). The results clearly illustrate that the conversational model, SkySenseGPT [58], maintains its exceptional performance across both tasks. Notably, none of the contrastive VLMs proved effective for the VQA task, further underscoring the advantages of conversational methods. Additionally, in the IC task, the prevalence of conversational VLMs exceeds that of their contrastive counterparts.
In conclusion, our analysis indicates that conversational VLMs not only demonstrate greater versatility across a range of tasks but also deliver superior performance overall. Consequently, we advocate for the adoption of conversational model architectures as the preferred choice for future developments in visual-language modeling.
6 Conclusions and Future Work
This review provides a systematic overview of Vision-Language Models (VLM) in the field of remote sensing. It introduces milestone works in the development of VLM itself, discusses the current research status of VLM in remote sensing, outlines the tasks involved, and offers a brief comparison of different VLM methods in various remote sensing applications. Overall, VLM has once again facilitated advancements in data processing technologies within the field of remote sensing. Current work has clearly demonstrated the potential of VLM applications in this area. However, there are still some shortcomings in current VLM technologies and their applications in remote sensing. Future work could consider the following directions:
  • Addressing regression problems. Currently, the mainstream approach in VLM involves using a tokenizer to vectorize text, then aligning visual information with the text word vectors before feeding them into the LLM model. This approach results in the model being insensitive to numerical values. For example, the number 100 may be tokenized as ’1’, ’00’, or it could be split into ’10’, ’0’, or ’100’. This leads to suboptimal performance when the current framework is applied to certain regression tasks. Future research on how to integrate regression problems into the current framework is a topic worth exploring.
  • Aligning with the structural characteristics of remote sensing images. Current VLM efforts in the remote sensing field largely follow the overarching framework established in the computer vision domain, without fully leveraging the unique characteristics of remote sensing images. For instance, remote sensing images come in various types, such as SAR images, HSI, and radar images, each with distinct features. Designing feature extractors that cater to these data characteristics is essential for maximizing the advantages of the images. However, in the current remote sensing VLM work, visual encoders primarily rely on pre-trained feature extractors based on RGB images, which do not fully align with remote sensing images. Additionally, there is a lack of research considering the fusion of different modalities of remote sensing images to enhance the performance of VLM.
  • Multimodal output. Existing VLM models in the remote sensing field typically output only text. While they can handle traditional tasks like image classification, they struggle with dense prediction tasks such as segmentation, where text output is insufficient. Therefore, researching multimodal outputs—allowing VLM models to also output images or even videos—is a highly valuable direction. This could facilitate the integration of various application needs within the remote sensing domain into a unified framework.